OpenStudy (anonymous):

Complete the ordered pairs so that each is a solution of the given linear equation. Then graph the equation. x=6+y Complete the ordered pair x and y (0, ) (-2, ) (6, )

OpenStudy (campbell_st):

ok... for the ordered pairs substitute x = 0 to find a matching y value. 0 = 6 + y ..... then y = -6 so the ordered pair is (0, -6) repeat the process for x = -2 and then x = 6 Once you have the 3 ordered pairs, plot then on the number plane. Draw a line through the 3 points.

OpenStudy (anonymous):

So, -2=6+y

OpenStudy (anonymous):

do I add -2 to 6

OpenStudy (campbell_st):

subtract 6 from both sides

OpenStudy (anonymous):

6 minus -2= 4

OpenStudy (anonymous):

(-2,4)

OpenStudy (campbell_st):

-2 = 6 + y -6 -6 ---------- -8 = y so you have an ordered pair of (-2, -8)

OpenStudy (anonymous):

so the third one is (6,0)

OpenStudy (anonymous):

wait 6=6+y -6 -6

OpenStudy (campbell_st):

thats it y = 0 so (6,0) just plot the 3 points on the number plane and you should be able to draw a straight line through all 3 points.

OpenStudy (anonymous):

ok